# GOT4Rec: Graph of Thoughts for Sequential Recommendation

### v1

@@ 3.5 Multi-Level Retrieval Module
Sequence-level retrieval. As previously mentioned, when analyzing collaborative preferences, it is essential to retrieve sequences of other users who share similar interests with the current user. We choose to deploy the all-mpnet-base-v2 model $f_{mpnet}$ due to its superior efficiency and discriminative power in encoding item titles compared to other encoding models such as BERT. Given a query sequence $S_{u}=\{i_{1},...,i_{k}\}$, the ranked indices of the retrieved sequences $I_{seq}$ can be obtained as follows:
$I_{seq}=sim(\frac{1}{k}\sum_{t=1}^{k}f_{mpnet}(i_{t}),V_{seq})$ (1)
where $sim(\cdot,\cdot)$ denotes the computation of Euclidean distance, $V_{seq}$ represents the vector base containing the embeddings of all other sequences.
Item-level retrieval. The title of an item generated by LLMs may differ from the title of the ground truth in datasets, even though they refer to the same item. This discrepancy arises due to the limited and varied versions of items in the datasets (e.g., Witcher 3: Wild Hunt versus Witcher 3: Wild Hunt Complete Edition). To address this issue, we continue to utilize the $f_{mpnet}$ model to encode item titles into vectors. We then retrieve the most similar items from the vector base by computing the inner product of the query vector and all other vectors in the base. Specifically, for an item $i_{query}$ generated by LLMs, the retrieval process is as follows:
$I_{item}=sim(f_{mpnet}(i_{query}),V_{item})$ (2)
where $I_{item}$ represents the ranked indices of the retrieved items, $sim(\cdot,\cdot)$ denotes the computation of the inner product and $V_{item}$ is the vector base containing all item embeddings. In practice, we observed that in most cases, when the description of the query item is vague, the retrieved items are misaligned (e.g., when querying for The Witcher 3, the top results include [Diablo III, The Witch and the Hundred Knight - PlayStation 3, The Witcher 3: Wild Hunt - Xbox One]). To mitigate this, we retrieve the top-$K$ items within the ranked indices for each recommended item.
@@ 4.1 Experimental Settings
Datasets. We conduct experiments on three item categories from the widely used Amazon Reviews’23 dataset (Hou et al. 2024a): Video Games (Games), Grocery and Gourmet Food (Food) and Home and Kitchen (Home). Reviews are treated as user-item interactions, sequenced chronologically by timestamp. To address the significant time cost of LLMs, we focus on users with 6 to 20 interactions and filter out items with fewer than five interactions. Following (Wang et al. 2024), we randomly sample 3,000 users from each dataset three times. We employ the leave-one-out strategy for dataset division: the most recent interaction for testing, the second most recent interaction for validation and the remaining interactions are used for training. Both training and validation segments are used as input during testing.
Dataset Metric GRU4Rec SRGNN SASRec SLIM CoT CoT-SC ToT GOT4Rec Improment
Games HR@5 0.0390 0.0312 0.0776 0.0602 0.0644 0.0653 0.0489 0.0894 15.21%
HR@10 0.0545 0.0495 0.0953 0.0977 0.0988 0.1013 0.0712 0.1167 15.20%
HR@20 0.0765 0.0817 0.1227 0.1281 0.1347 0.1253 0.1087 0.1361 1.04%
NDCG@5 0.0266 0.0220 0.0528 0.0380 0.0408 0.0421 0.0304 0.0621 17.61%
NDCG@10 0.0317 0.0277 0.0586 0.0502 0.0519 0.0537 0.0377 0.0710 21.16%
NDCG@20 0.0372 0.0357 0.0655 0.0578 0.0609 0.0622 0.0471 0.0760 16.03%
Food HR@5 0.0213 0.0307 0.0335 0.0350 0.0396 0.0443 0.0273 0.0742 67.49%
HR@10 0.0329 0.0352 0.0407 0.0517 0.0581 0.0597 0.0390 0.0972 62.81%
HR@20 0.0525 0.0528 0.0549 0.0613 0.0748 0.0753 0.0533 0.1090 44.75%
NDCG@5 0.0133 0.0196 0.0286 0.0216 0.0253 0.0276 0.0182 0.0492 72.03%
NDCG@10 0.0170 0.0243 0.0309 0.0270 0.0311 0.0326 0.0219 0.0567 73.93%
NDCG@20 0.0219 0.0288 0.0337 0.0295 0.0352 0.0366 0.0255 0.0597 63.11%
Home HR@5 0.0095 0.0068 0.0132 0.0123 0.0118 0.0133 0.0047 0.0192 44.36%
HR@10 0.0164 0.0115 0.0179 0.0180 0.0177 0.0223 0.0100 0.0299 34.08%
HR@20 0.0274 0.0197 0.0262 0.0213 0.0230 0.0270 0.0147 0.0337 24.81%
NDCG@5 0.0058 0.0049 0.0097 0.0073 0.0073 0.0080 0.0031 0.0122 25.77%
NDCG@10 0.0080 0.0065 0.0105 0.0091 0.0093 0.0109 0.0049 0.0157 44.04%
NDCG@20 0.0108 0.0085 0.0134 0.0099 0.0106 0.0121 0.0060 0.0167 24.63%
@@
Baselines. To demonstrate the effectiveness of our proposed method, we select two groups of recommendation baselines. The first group comprises traditional neural sequential recommendation models:
  • GRU4Rec (Hidasi et al. 2016): Uses RNNs to model user sequences for session-based recommendation.
  • SRGNN (Wu et al. 2019): A graph-based model for session-based recommendation to capture the transition information between items in user sequences.
  • SASRec (Kang and McAuley 2018): A self-attention based sequential model to capture user’s preferences.
The second group contains models that utilize LLM prompting strategies:
  • Chain-of-Thought (CoT) (Wei et al. 2022): A prompting approach for prompting which includes the intermediate steps of reasoning within the prompt.
  • Multiple CoTs (CoT-SC) (Wang et al. 2023): A scheme in which multiple CoTs are generated, with the best one being selected as final result.
  • Tree of Thoughts (ToT) (Yao et al. 2024): A prompting approach modeling the LLM reasoning process as a tree.
  • SLIM (Wang et al. 2024): A knowledge distillation module, which transfers the step-by-step reasoning capabilities in recommendation from a larger teacher model to a smaller student model.
Implementation Details. We choose Llama3-8B-Instruct (AI@Meta 2024) as the backbone model with a maximum token length of 4096. The LMDeploy toolkit (Contributors 2023) is employed to deploy LLMs and accelerate reasoning. GRU4Rec, SASRec and SRGNN are implemented based on their official code, and LLM prompting strategies are integrated within the GoT framework. For SLIM, Llama3-8B-Instruct is used instead of ChatGPT to avoid high API costs. We retrieve the 10 most similar items for both GOT4Rec and baseline methods using the faiss library (Douze et al. 2024). Optimal hyper-parameters for all baselines are carefully selected to ensure the best performance.
Evaluation Metrics. We evaluate performance using hit rate (HR) and normalized discounted cumulative gain (NDCG), reporting HR@K and NDCG@K for $K\in\{5,10,20\}$. Each recommended item is evaluated against all other items in the sampled datasets. To assess recommendation novelty, we calculate EFD@10 and EPC@10 (Vargas and Castells 2011). The average scores of three runs are reported.
@@ 4.2 Overall Performance
Table 1 compares our GOT4Rec method with various neural sequential models and LLM prompting strategies, leading to several key observations: (1) Traditional neural sequential models perform relatively modest, though SASRec still outperforms LLM prompting strategies in some cases, particularly due to its use of self-attention mechanisms which allow SASRec to effectively model users’ historical behaviors by capturing the transition relationships between items. However, SASRec lacks the ability to comprehend semantic information, limiting its overall performance. (2) Among LLM-based methods, CoT-SC consistently achieves runner-up performance across most datasets, largely because it aggregates and selects the best result from multiple CoT paths, providing a more refined output. On the other hand, CoT, ToT, and SLIM show comparatively lower performance. SLIM, in particular, may suffer from reduced output diversity due to fine-tuning, while ToT’s structural and reasoning path designs appear to be less suitable for sequential recommendation tasks. (3) GOT4Rec achieves the state-of-art (SOTA) performances across all datasets. Notably, in the Food dataset, GOT4Rec achieves a relative improvement of 73.93% over CoT-SC in terms of NDCG@10 and 67.49% in terms of HR@5. These significant gains can be attributed to the nature of food product consumption, where users prefer consistent categories or brands and are strongly influenced by short-term needs. This aligns well with the strengths of GOT4Rec, which excels at capturing users’ preferences for certain categories or brands and effectively integrating short-term preference information. This capability allows GOT4Rec to deliver highly relevant recommendations that closely match users’ current interests. These findings demonstrate the effectiveness of GOT4Rec in optimizing recommendation tasks by fully exploiting the advanced reasoning capabilities of LLMs and integrating various aspects of user information.
Dataset Metric $w/o$ Short. $w/o$ Long. $w/o$ Co. GOT4Rec
Games HR@5 0.0819 0.0871 0.0774 0.0894
HR@10 0.1083 0.1121 0.1015 0.1167
HR@20 0.1262 0.1337 0.1170 0.1361
NDCG@5 0.0512 0.0573 0.0527 0.0621
NDCG@10 0.0597 0.0654 0.0606 0.0710
NDCG@20 0.0642 0.0709 0.0645 0.0760
Food HR@5 0.0591 0.0867 0.0687 0.0742
HR@10 0.0867 0.0933 0.0880 0.0972
HR@20 0.1003 0.1063 0.1013 0.1090
NDCG@5 0.0362 0.0426 0.0437 0.0492
NDCG@10 0.0453 0.0509 0.0500 0.0567
NDCG@20 0.0487 0.0543 0.0534 0.0597
Home HR@5 0.0179 0.0166 0.0190 0.0192
HR@10 0.0284 0.0292 0.0270 0.0299
HR@20 0.0322 0.0318 0.0309 0.0337
NDCG@5 0.0102 0.0103 0.0114 0.0122
NDCG@10 0.0136 0.0144 0.0140 0.0157
NDCG@20 0.0146 0.0150 0.0149 0.0167
@@
Table 2: Ablation analysis, conducted by retaining different components in GOT4Rec to form variants. The best performance is highlighted in bold. “Short.”, “Long.”, and “Co.” denote long-term, short-term, and collaborative schemes, respectively.
@@ 4.3 Ablation Study
We conducted an ablation study to analyze the impact of different components in the GOT4Rec model, with the results in Table 2. Our GOT4Rec consistently outperforms the ablated variants, demonstrating that the full integration of users’ preference information from the short-term, long-term, and collaborative components results in superior recommendation performance. The study also reveals that the importance of each component varies across different datasets, likely reflecting the unique characteristics of each dataset. For instance, collaborative information appears to be the most crucial component in Games dataset. This suggests that when purchasing gaming products, users prioritize categories or brands, making collaborative information from other users with similar interests particularly important. In contrast, short-term preference plays a more significant role in Food dataset, aligning with our inference that users’ recent preferences heavily influence their choices in food products. In Home dataset, the impact of each component varies, but short-term preference has the least influence. This could indicate that long-term preferences or collaborative insights are more critical when users make decisions about home-related items.
Figure 4: Analysis of popularity bias in Games dataset. Items in the dataset are sorted by their frequency. Compared to CoT, GOT4Rec demonstrates a more consistent ability to recommend long-tail items.
@@ 4.4 Popularity Bias Analysis
In Figure 4, we sort the items in Games dataset based on their frequency in the training set (i.e., popularity) and draw lines to illustrate each item’s frequency in the results of CoT and GOT4Rec. The figures for other two datasets are presented in the Appendix. It is apparent that GOT4Rec more effectively recommends tail items and a broader variety of items. Table 3 reports EFD@10 and EPC@10 metrics for CoT, SLIM and GOT4Rec, indicating tthat GOT4Rec outperforms the baselines in recommending long-tail items. These results support our claim that GOT4Rec mitigates popularity bias by capturing a wider range of information.
@@
Table 3: Popularity bias metrics, the higher score indicates higher recommendation novelty. The best performance is highlighted in bold.

### v2

@@ 3.4 Multi-Level Retrieval Module
Sequence-level Retrieval.
As previously mentioned, when analyzing collaborative preferences, it is essential to retrieve sequences of other users who share similar interests with the current user. We choose to deploy the Mpnet Reimers and Gurevych (2020) model $f_{\text{mpnet}}$ due to its superior efficiency and discriminative power in encoding item titles compared to other encoding models such as BERT. Given a query sequence $S_{u}=\{i_{1},...,i_{m}\}$, the retrieved sequences set ${\mathcal{S}_{\mathcal{V}_{\text{seq}}}^{k}}$ can be obtained as follows:
${\mathcal{S}_{\mathcal{V}_{\text{seq}}}^{k}}(i_{t})\triangleq{\arg\max_{\boldsymbol{v}\in\mathcal{V}_{\text{seq}}}^{k}}sim(\frac{1}{m}\sum_{t=1}^{m}f_{\text{mpnet}}(i_{t}),\boldsymbol{v})$ (1)
where ${\mathcal{S}_{\mathcal{V}_{\text{seq}}}^{k}}$ is the set of top-$k$ sequences that share similar interests with the current user. $sim(\cdot,\cdot)$ denotes the computation of Euclidean distance, $V_{\text{seq}}$ represents the vector base containing the embeddings of all other sequences.
Item-level Retrieval.
The title of an item generated by LLMs may differ from the title of the ground truth in datasets, even though they refer to the same item. This discrepancy arises due to the limited and varied versions of items in the datasets (e.g., Witcher 3: Wild Hunt versus Witcher 3: Wild Hunt Complete Edition). To address this issue, we continue to utilize the $f_{mpnet}$ model to encode item titles into vectors. We then retrieve the most similar items from the vector base by computing the inner product of the query vector and all other vectors in the base. Specifically, for an item $i_{\text{query}}$ generated by LLMs, the retrieval process is as follows:
${\mathcal{I}_{\mathcal{V}_{\text{item}}}^{k}}(i_{\text{query}})\triangleq{\arg\max_{\boldsymbol{v}\in\mathcal{V}_{\text{item}}}^{k}}sim(f_{\text{mpnet}}(i_{\text{query}}),\boldsymbol{v})$ (2)
where ${\mathcal{I}_{\mathcal{V}_{\text{item}}}^{k}}$ represents the set of the retrieved top-$k$ items based on $sim(\cdot,\cdot)$, which denotes the computation of the inner product and $V_{\text{item}}$ is the vector base containing all item embeddings. In practice, we observed that in most cases, when the description of the query item is vague, the retrieved items are misaligned (e.g., when querying for The Witcher 3, the top results include [Diablo III, The Witch and the Hundred Knight - PlayStation 3, The Witcher 3: Wild Hunt - Xbox One]). To mitigate this, we retrieve the top-$K$ items within the ranked indices for each recommended item.
3.5 Inference Latency and Volume Analyzation
In Table 1, we compare the inference lantency and volume between reasoning frameworks. Lantency is the number of hops to reach the final thought and volume refer to the number of preceding thoughts that have impacted the final thought Besta et al. (2024). The time to generate a single thought is assumed to be $O(1)$. In CoT, reasoning follows a single sequential chain, resulting in both latency and volume scaling with $N$. CoT-SC employs $k$ independent chains, noting that these chains support parallel execution, both are reduced to $N/k$. For a $k$-ray ToT tree, both latency and volume are $\log_{k}N$. As for GoT, it is a $k$-ray tree combined with its inverted counterpart, resulting in a latency of $\log_{k}N$. For volume, since all intermediate thoughts contribute to the final thought by aggregating information from every step, it scales with $N$.
Framework Latency Volume
Chain-of-Thought (CoT) $N$ $N$
Multiple CoTs (CoT-SC) $N/k$ $N/k$
Tree of Thoughts (ToT) $\log_{k}N$ $\log_{k}N$
Graph of Thoughts (GoT) $\log_{k}N$ $N$
Table 1: Comparison of inference latency and volume for different reasoning frameworks.
@@ 4.1 Experimental Settings
Datasets.
We conduct experiments on three item categories from Amazon Reviews’23 dataset Hou et al. (2024a): Video Games (Games), Grocery and Gourmet Food (Food) and Home and Kitchen (Home). Reviews are treated as user-item interactions, sequenced chronologically by timestamps. We focus on users with 6 to 20 interactions and filter out items with fewer than five interactions. Following Wang et al. (2024), we randomly sample 3,000 users from each dataset three times. We employ the leave-one-out strategy for dataset division: the most recent interaction for testing, the second most recent interaction for validation and the remaining are used for training. Both training and validation segments are used as input during testing.
Dataset Metric SASRec PALR TALLRec SLIM CoT CoT-SC ToT GOT4Rec Improv.
Games HR@5 0.0776 0.0630 0.0660 0.0602 0.0644 0.0653 0.0489 0.0894 15.21%
HR@10 0.0953 0.0893 0.0967 0.0977 0.0988 0.1013 0.0712 0.1167 15.20%
HR@20 0.1227 0.1157 0.1233 0.1281 0.1347 0.1253 0.1087 0.1361 1.04%
NDCG@5 0.0528 0.0385 0.0401 0.0380 0.0408 0.0421 0.0304 0.0621 17.61%
NDCG@10 0.0586 0.0423 0.0497 0.0502 0.0519 0.0537 0.0377 0.0710 21.16%
NDCG@20 0.0655 0.0475 0.0562 0.0578 0.0609 0.0622 0.0471 0.0760 16.03%
Food HR@5 0.0335 0.0337 0.0390 0.0350 0.0396 0.0443 0.0273 0.0742 67.49%
HR@10 0.0407 0.0487 0.0513 0.0517 0.0581 0.0597 0.0390 0.0972 62.81%
HR@20 0.0549 0.0657 0.0747 0.0613 0.0748 0.0753 0.0533 0.1090 44.75%
NDCG@5 0.0286 0.0216 0.0243 0.0216 0.0253 0.0276 0.0182 0.0492 72.03%
NDCG@10 0.0309 0.0235 0.0284 0.0270 0.0311 0.0326 0.0219 0.0567 73.93%
NDCG@20 0.0337 0.0267 0.0318 0.0295 0.0352 0.0366 0.0255 0.0597 63.11%
Home HR@5 0.0132 0.0130 0.0123 0.0123 0.0118 0.0133 0.0047 0.0192 44.36%
HR@10 0.0179 0.0167 0.0183 0.0180 0.0177 0.0223 0.0100 0.0299 34.08%
HR@20 0.0262 0.0213 0.0233 0.0213 0.0230 0.0270 0.0147 0.0337 24.81%
NDCG@5 0.0097 0.0080 0.0081 0.0073 0.0073 0.0080 0.0031 0.0122 25.77%
NDCG@10 0.0105 0.0094 0.0096 0.0091 0.0093 0.0109 0.0049 0.0157 44.04%
NDCG@20 0.0134 0.0101 0.0113 0.0099 0.0106 0.0121 0.0060 0.0167 24.63%
@@
Baselines.
To demonstrate the effectiveness of our proposed method, we select two groups of recommendation baselines. The first group comprises both neural and LLM-based sequential recommenders:
  • SASRec Kang and McAuley (2018): A self-attention based sequential model to capture user’s preferences.
  • PALR Yang et al. (2023): A novel framework that integrates user behavior with LLMs to generate personalized recommendations.
  • TALLRec Bao et al. (2023): An efficient fine-tuning framework that aligns LLMs with recommendation systems through a two-stage tuning process.
The second group contains models that utilize LLM reasoning strategies:
  • SLIM Wang et al. (2024): A knowledge distillation module transferring the step-by-step reasoning capabilities in recommendation from a larger teacher model to a smaller student model.
  • Chain-of-Thought (CoT) Wei et al. (2022): A reasoning approach for prompting which includes the intermediate steps of reasoning within the prompt.
  • Multiple CoTs (CoT-SC) Wang et al. (2023): A scheme in which multiple CoTs are generated, with the best one being selected as final result.
  • Tree of Thoughts (ToT) Yao et al. (2024): A reasoning approach modeling LLM reasoning process as a tree.
Implementation Details.
We choose Llama3-8B-Instruct AI@Meta (2024) as the backbone model. SASRec and TALLRec are implemented based on their official code. We implemented PALR by ourselves and LLM reasoning strategies are integrated within the GoT framework. For SLIM, Llama3-8B-Instruct is used instead of ChatGPT to avoid high API costs. We retrieve the 10 most similar items for both GOT4Rec and baseline methods using the faiss library Douze et al. (2024). Optimal hyper-parameters for all baselines are carefully selected to ensure the best performance.
Evaluation Metrics.
We evaluate performance with hit rate (HR) and normalized discounted cumulative gain (NDCG), reporting HR@K and NDCG@K for $K\in\{5,10,20\}$. Each recommended item is evaluated against all other items in the dataset. The average scores of three runs are reported.
@@ 4.2 Overall Performance
Table 2 compares our GOT4Rec method with various neural sequential models and LLM reasoning strategies, leading to several key observations:
Performance of Neural & LLM-based Models.
Neural and LLM-based sequential models perform relatively modest, though SASRec still outperforms LLM reasoning strategies in some cases, particularly due to its use of self-attention mechanisms which allow SASRec to effectively model users’ historical behaviors by capturing the transition relationships between items. However, SASRec lacks the ability to comprehend semantic information, limiting its overall performance. As for PALR and TALLRec, their suboptimal performance stems from the absence of strong reasoning capabilities, which are critical for complex recommendation tasks.
Performance of LLM Reasoning Methods.
Among LLM reasoning methods, CoT-SC consistently achieves runner-up performance across most datasets, largely because it aggregates and selects the best result from multiple CoT paths, providing a more refined output. On the other hand, CoT, ToT, and SLIM show comparatively lower performance. SLIM, in particular, may suffer from reduced output diversity due to fine-tuning, while ToT’s structural and reasoning path designs appear to be less suitable for sequential recommendation tasks.
Superiority of GOT4Rec.
@@
GOT4Rec achieves the state-of-art performances across all datasets. Notably, in the Food dataset, GOT4Rec achieves a relative improvement of 73.93% over CoT-SC in terms of NDCG@10 and 67.49% in terms of HR@5. These significant gains can be attributed to the nature of food product consumption, where users prefer consistent categories or brands and are strongly influenced by short-term needs. This aligns well with the strengths of GOT4Rec, which excels at capturing users’ preferences for certain categories or brands and effectively integrating short-term preference information. This capability allows GOT4Rec to deliver highly relevant recommendations that closely match users’ interests. These findings demonstrate the effectiveness of GOT4Rec in optimizing recommendation tasks by fully exploiting the advanced reasoning capabilities of LLMs and integrating various aspects of user information.
Dataset Metric $w/o$ Short-term $w/o$ Long-term $w/o$ Collaborative GOT4Rec
Games HR@5 0.0819↓8.39 0.0871↓2.57 0.0774↓13.42 0.0894
HR@10 0.1083↓7.20 0.1121↓3.94 0.1015↓13.02 0.1167
HR@20 0.1262↓7.27 0.1337↓1.76 0.1170↓14.03 0.1361
NDCG@5 0.0512↓17.55 0.0573↓7.73 0.0527↓15.14 0.0621
NDCG@10 0.0597↓15.92 0.0654↓7.89 0.0606↓14.65 0.0710
NDCG@20 0.0642↓13.53 0.0709↓6.71 0.0645↓15.13 0.0760
Food HR@5 0.0591↓20.35 0.0867↑16.85 0.0687↓7.41 0.0742
HR@10 0.0867↓10.80 0.0933↓4.01 0.0880↓9.47 0.0972
HR@20 0.1003↓7.98 0.1063↓2.48 0.1013↓7.06 0.1090
NDCG@5 0.0362↓26.42 0.0426↓13.41 0.0437↓11.18 0.0492
NDCG@10 0.0453↓20.11 0.0509↓10.23 0.0500↓11.82 0.0567
NDCG@20 0.0487↓18.43 0.0543↓9.05 0.0534↓10.55 0.0597
Home HR@5 0.0179↓6.77 0.0166↓13.54 0.0190↓1.04 0.0192
HR@10 0.0284↓5.02 0.0292↓2.34 0.0270↓9.70 0.0299
HR@20 0.0322↓4.45 0.0318↓5.64 0.0309↓8.31 0.0337
NDCG@5 0.0102↓16.39 0.0103↓15.57 0.0114↓6.56 0.0122
NDCG@10 0.0136↓13.38 0.0144↓8.28 0.0140↓10.83 0.0157
NDCG@20 0.0146↓12.57 0.0150↓10.18 0.0149↓10.78 0.0167
@@
Table 3: Ablation analysis, conducted by retaining different components in GOT4Rec to form variants. The best performance is highlighted in bold. The performance difference (%) with GOT4Rec is highlighted with blue and orange. “Short-term”, “Long-term”, and “Collaborative” denote short-term, long-term and collaborative reasoning steps, respectively.
@@ 4.3 Ablation Study
We conducted an ablation study to analyze the impact of different components in the GOT4Rec model, with the results in Table 3. Our GOT4Rec consistently outperforms the ablated variants, demonstrating that the full integration of users’ preference information from the short-term, long-term, and collaborative components results in superior recommendation performance. The study also reveals that the importance of each component varies across different datasets, likely reflecting the unique characteristics of each dataset. For instance, collaborative information appears to be the most crucial component in Games dataset. This suggests that when purchasing gaming products, users prioritize categories or brands, making collaborative information from other users with similar interests particularly important. In contrast, short-term preference plays a more significant role in the Food dataset, as users’ recent interactions are often influenced by immediate needs, cravings, or specific dietary goals. This aligns with our inference that users’ recent preferences heavily influence their choices in food products. In the Home dataset, the impact of each component varies, but short-term preference has the least influence. This indicates that long-term preferences or collaborative insights are more critical when users make decisions about home-related items that typically involve more thoughtful consideration.
@@ 4.4 Popularity Bias Analysis
To assess recommendation novelty, we calculate EFD@10 and EPC@10 Vargas and Castells (2011). EFD (Expected Free Discovery) represents the expected inverse collection frequency of relevant and seen recommended items. EPC (Expected Popularity Complement) measures the expected number of seen relevant recommended items that were previously unknown to the user. These two metrics provide a comprehensive assessment of diversity and novelty, helping to balance the recommendation of popular items with long-tail content. Table 4 reports EFD@10 and EPC@10 metrics for CoT, SLIM and GOT4Rec, indicating that GOT4Rec outperforms the baselines in recommending long-tail items. These results support our claim that GOT4Rec mitigates popularity bias by capturing a wider range of information.
@@
Table 4: Popularity bias metrics, the higher score indicates higher recommendation novelty. The best score is highlighted in bold.